May 20, 2014 Testing

Goal:

  • Figure out weird frequency switch issue: we're getting lines where we shouldn't in the switch

Plan:

  • inject test tone in mode 15 with no switching and see if we see the problem
  • turn on frequency switching in mode 15 and see if we see a problem
  • Also try mode 1. that may help isolate issues.

Info:

  • TGBT13B_502, session 62

Log:

  • Scan 1: fsw_issue_noswitch_nocal. mode 15, one window. test tone at 1421.405. aborted. tone not set up properly
  • Scan 2: fsw_issue_noswitch_nocal. aborted. still trying to get tone set up properly.
  • Scan 3: fsw_issue_noswitch_nocal. added some code to set up test tone automatically.
  • Scan 4: fsw_issue_noswitch_nocal_notone. looking at it with no tone.
  • Scan 5: fsw_issue_noswitch_nocal_notone. switched to topo.
  • Scan 6: fsw_issue_noswitch_nocal. switched to topo. VEGAS is consistently taking 30s beyond the end of a scan to finish.
  • Scan 7: fsw_issue_noswitch_nodcal_notone_mode1. checking for extended running in mode 1. finished immediately.
  • Scan 8: fsw_issue_noswitch_nocal. running so that Joe can see what happens at the end of a scan. see tone, but also see signal at freq=1415.55 MHz, chan=29886. seen in scan without tone.
  • Scan 9: fsw_issue_noswitch_nocal. moved window to 1419.405 from 1420.405 to see how the tone and the signal move. possible image of HI at edge of spectrum? Don't see in earlier scans.
  • Scan 10: fsw_issue_noswitch_nocal. moved window to 1418.405 to see if possible image moves.
  • Scan 11: fsw_issue_noswitch_nocal_mode16. Changed to mode 16 to see if still have the same issue. center at 1420.405.
  • Scan 12: fsw_issue_noswitch_nocal_mode16. center at 1419.405
  • Scan 13: fsw_issue_noswitch_nocal_mode16. center at 1418.405.
  • Joe switched versions of hpc software to old version
  • Scan 14: fsw_issue_noswitch_nocal_mode 16. center at 1418.405. vegas faulted.
  • wrong version of hpc software. Joe fixed.
  • Scan 15: fsw_issue_noswitch_nocal_mode 16. center at 1418.405. same behavior with the old and new versions. images at feature-1/2*bw
  • Scan 16: fsw_issue_noswitch_nocal_mode7. Run in non-tunable LO mode to see if that's an issue. center at 1418.405. not seeing same images.
  • Scan 17: fsw_issue_noswitch_nocal_mode7. center 1410.405 to make sure we don't see a test tone image. don't see test tone image.
  • Scan 18: fsw_issue_noswitch_nocal_mode7. center 1440.405 to make sure that the test tone on the other side. nope. test tone isn't on other side.
  • Scan 19: fsw_issue_noswitch_nocal_mode10. center 1418.405. see images in mode 10.
  • Scan 20: fsw_issue_noswitch_nocal_mode20. center 1418.405. no image.
  • Scan 21: fsw_issue_noswitch_nocal_mode25. center 1418.405. no image.
  • lunch
  • moving window around manually
  • Scan 22: fsw_issue_noswitch_nocal. Set subfrequencyA,1 to 249000000 Hz to see how changing the window location without changing the LO works. image moved to 11.72+1/2*bw
  • Scan 23: fsw_issue_noswitch_nocal. Set subfrequencyA,1 to 251000000 Hz. image at 11.72-1/2*bw. suggests repeating image.
  • Scan 24: fsw_issue_noswitch_nocal_mode4.
  • Ray updated manager to check for zeros at GPU.
  • Scan 25: fsw_issue_mode4. Trying frequency switching with tone in modes without the image problem to see what works. saw zero channel issue. Ray did not receive any error messages from GPU code about zeros. UPDATE: He IS seeing error messages about zeros. Frequency switching looks fine.
  • Scan 26: fsw_issue_mode7. looks okay.
  • Scan 27: fsw_issue_mode10. This should fail. It didn't fail, but the data look really weird.
  • Scan 28: fsw_issue_mode10. Trying again to see if can replicate the problem. Ugh. Doesn't look good. Lots of zeros. Going through and checking total power and then switched power.
  • Scan 29: fsw_issue_noswitch_nocal_mode10. Checking total power. Still ugly
  • restarting HPC software
  • Scan 30: fsw_issue_noswitch_nocal_mode10. This looks okay.
  • Scan 31-34: fsw_issue_noswitch_nocal_mode10. Queueing up a bunch to run. While Ray is looking at HPC software.
  • Scan 35: fsw_issue_mode20.
  • Scan 36: fsw_issue_mode25.
  • Scan 37: fsw_issue_mode25. Shifting center to see if there's images in both.
  • Scan 38: fsw_issue_mode16. Running to check individual integrations.
  • Scan 39: fsw_issue_noswitch_nocal_mode16. Changed lbw_gain,1 from 1024 to 2048 at Glenn's suggestion.
  • Scan 40: fsw_issue_noswitch_nocal_mode16. Changed lbw_gain,1 to 256 to see what happens
  • Scan 41: fsw_issue_noswitch_nocal_mode15. Changed tint=1s to see if that helps. It does! Now what happens with the frequency switch. UPDATE: No, it doesn't. Just couldn't see HI.
  • Scan 42: fsw_issue_mode15. Changed tint=1s and frequency switch is still bad.

-- AmandaKepley - 2014-05-20
Topic revision: r3 - 2015-01-27, RichardPrestage
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding NRAO Public Wiki? Send feedback